40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 14 01:19:18 2011, rana, Update, CDS, Today's CDS problems 
    Reply  Mon Feb 14 07:58:43 2011, Suresh, Update, CDS, Today's CDS problems 
Message ID: 4282     Entry time: Mon Feb 14 01:19:18 2011     Reply to this: 4285
Author: rana 
Type: Update 
Category: CDS 
Subject: Today's CDS problems 

This is just a listing of  CDS problems I still notice today:

  1. MC2-MCL button was left ON due to BURT failure. This, of course, screws up our Green locking investigations because of the unintended feedback. Please fix the BURT/button issue.

  2. The GCV - FB0 status is RED. I guess this means there's something wrong? Its really a bad idea to have a bunch of whited out or falsely red indicators. No one will ever use these or trust these in the future.
  3. MC1/2/3 Lockins are all white. Also, the MODE switches for the dewhitening are all white.
  4. Is the MC SIDE coil dewhitening filter synced with anything? It doesn't seem to switch anything. Maybe the dewhite indicators at the top right of the SUS screens can be made to show the state of the binary output instead of just the digital filter
  5. MC WFS is all still broken. We need a volunteer to take this on - align beams, replace diodes, fix code/screens.
ELOG V3.1.3-